# Gravastar on the brane with a timelike extra dimension

### v1

@@ Abstract
We construct a gravastar configuration within the Shtanov–Sahni (SS) braneworld scenario, characterized by a timelike extra dimension and negative brane tension. Unlike classical black holes, which inevitably culminate in central curvature singularities, our model demonstrates that the SS braneworld dynamics naturally regularize the interior geometry and prevent singularity formation. By solving the modified Einstein field equations induced on the brane, we obtain explicit interior, shell, and exterior solutions without invoking the idealized thin-shell approximation. The gravastar core is modeled as a Bose–Einstein condensate, while the intermediate shell consists of ultra-dense stiff matter. Bulk Weyl corrections induce anisotropic effective pressures on the brane, a feature that emerges intrinsically in this scenario and supports stability. We analyze the active gravitational mass, energy, entropy, and proper thickness of the shell, and establish the junction conditions at the interfaces. Our analysis reveals that the SS gravastar exhibits suppressed or even negative effective mass, reflecting the repulsive nature of the interior condensate, and admits stable equilibrium solutions consistent with energy conditions. This highlights the SS braneworld gravastar as a physically viable compact object and a compelling alternative to black holes. A key novelty of our construction is that the stabilizing pressure anisotropy and suppressed effective gravitational mass arise dynamically from higher-dimensional Weyl corrections, rather than being imposed through ad hoc matter sources or thin-shell idealizations. This provides the first fully analytic realization of a finite-thickness, stable gravastar in the Shtanov–Sahni braneworld, highlighting a genuinely geometric mechanism for singularity avoidance in compact objects.
1 Introduction
@@
The question of how gravitational collapse ultimately ends in astrophysics mirrors, in many respects, the problem of the initial state of the Universe in cosmology. While Einstein’s General Relativity (GR) has enjoyed remarkable success in explaining a wide range of observational phenomena at intermediate energy scales in both astrophysical and cosmological contexts, its predictive power is severely challenged in regimes of extreme curvature and energy density. Such regimes naturally arise at the putative beginning of the Universe and at the final stages of stellar gravitational collapse. In these limits, enormous amounts of energy are compressed into extremely small volumes, causing the energy density and curvature invariants to diverge. As a consequence, GR generically predicts spacetime singularities where the classical field equations lose their validity [1]. This pathological behavior strongly suggests that new physical effects, most plausibly of quantum origin, must become relevant in order to regularize the theory and remove these unphysical infinities.
@@
In the context of stellar collapse, incorporating quantum field effects on a fixed classical spacetime background has already revealed several profound phenomena. The most celebrated example is Hawking radiation, whereby black holes emit thermal radiation from their event horizons [2], often interpreted as the boundary separating the black hole interior from the external universe. Despite this remarkable insight, semiclassical gravity is insufficient to cure the singular behavior of classical black hole solutions, as the curvature singularity at the center persists unchanged [3]. In the Schwarzschild solution, the apparent singularity at the Schwarzschild radius R=2GM(c=1)R=2GM~(c=1) is merely a coordinate artifact, since all curvature invariants remain finite there and the singularity can be eliminated by an appropriate change of coordinates. In contrast, the divergence at $r=0$ corresponds to a genuine physical singularity, characterized by divergent curvature scalars, and cannot be removed by any coordinate transformation.
@@
From a fundamental perspective, the emergence of physical singularities in the description of any physical system is widely regarded as signaling the breakdown or incompleteness of the underlying theoretical framework. Relativistic gravity presents two prominent and closely related examples of such singularities: those associated with cosmological models of the Universe [4] and those arising in black hole spacetimes [5]. Both are curvature singularities, meaning that they are invariantly defined through divergences of curvature tensors and cannot be eliminated by coordinate redefinitions [6]. In standard relativistic cosmology, a Ricci-type curvature singularity occurs at $t=0$, where the Riemann tensor and the energy density diverge, marking what is conventionally interpreted as the beginning of spacetime evolution [4]. Analogously, classical black hole solutions contain a curvature singularity at $r=0$, where physical quantities blow up in an identical manner [5]. This central singularity stands in sharp contrast to the horizon at $r=2M$, which represents only a removable coordinate singularity in Schwarzschild coordinates and can be regularized through alternative coordinate systems such as those introduced by Finkelstein, Kruskal, and Szekeres [7, 8, 9].
In modern theoretical physics, it is increasingly accepted that the presence of such singularities reflects the limited domain of applicability of classical GR rather than a true physical feature of nature. This viewpoint is motivated by two key considerations. First, GR has not been experimentally tested in regimes of ultra-high curvature or energy density. Second, it is natural to expect that quantum effects should dominate the dynamics at such extreme scales, whereas GR is intrinsically a classical theory. Consequently, substantial effort has been devoted to extending or modifying GR in a manner that remains consistent with known observations while improving its high-energy behavior.
@@
One particularly influential strategy is to modify the geometric sector of the theory by generalizing the Einstein–Hilbert (EH) action. Originally introduced several decades ago as a means of incorporating ultraviolet (UV) corrections, such modifications have since been widely explored at both UV and infrared (IR) scales. This approach is often favored because it avoids many of the conceptual and phenomenological difficulties that arise when the matter sector alone is altered. Moreover, modifying the gravitational action may provide valuable clues toward the eventual construction of a self-consistent theory of quantum gravity (QG), even though such a complete theory remains elusive at present. In this regard, the early Universe and compact objects like black holes constitute natural laboratories for probing the viability of QG-inspired models, as both are characterized by extreme curvature and energy density.
Among the various approaches to quantum gravity, two broad and conceptually distinct frameworks have gained widespread attention. One of these is M-theory, which emerged as a unifying framework linking the five previously known superstring theories through duality relations and which reduces to supergravity in the low-energy limit [10, 11]. Quantum consistency of M-theory requires an eleven-dimensional spacetime, and its fundamental degrees of freedom include extended objects such as two- and five-dimensional membranes, commonly referred to as branes. Although M-theory is not yet fully formulated, its structural features have inspired numerous developments in high-energy physics and cosmology.
Braneworld models have emerged as an important extension of higher-dimensional gravity, with the Randall–Sundrum (RS) construction proposed in 1999 [12, 13] playing a central role. This scenario provides an elegant explanation of the hierarchy between the gravitational and electroweak scales [14] by postulating that our observable Universe is a $(3+1)$-dimensional hypersurface embedded in a higher-dimensional spacetime. In this setup, Standard Model fields are confined to the brane, whereas gravity propagates in the surrounding five-dimensional bulk endowed with an Anti–de Sitter geometry ($AdS_{5}$). The presence of a negative bulk cosmological constant leads to the localization of the graviton near the brane. While the original RS proposal involved two branes, the later RS-II formulation [13] effectively removes the negative-tension brane by sending it to infinity, yielding a single-brane configuration.
@@
The Randall–Sundrum single–brane framework with a spacelike extra dimension has been widely investigated and applied to a variety of problems in cosmology [15, 16, 17, 18, 19, 20, 21, 22, 23] as well as in astrophysical contexts [24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Building on this line of research, Shtanov and Sahni proposed an alternative construction, commonly referred to as the SS model, which may be regarded as a dual of the RS–II scenario, where the spacelike extra dimension is replaced by a timelike one [34]. In the standard RS case the five–dimensional bulk possesses a Lorentzian signature, whereas introducing a timelike extra dimension leads to a bulk geometry with signature (,,+,+,+)(-,-,+,+,+). This change in the nature of the extra dimension significantly modifies the properties of the braneworld. Unlike the RS–II model, the SS setup is characterized by a negative brane tension and a positive bulk cosmological constant. One of its most remarkable features is that the cosmological evolution admits a smooth bounce during the contracting phase, thereby avoiding the formation of a singularity, even when the matter content on the brane satisfies the usual energy conditions. A potential concern in models with a timelike extra dimension is the emergence of tachyonic Kaluza–Klein gravitational modes, as originally discussed in [34]; this issue has, however, been subsequently examined and addressed in [35].
Against this broader backdrop, Mazur and Mottola (MM) proposed in 2001 a novel alternative to classical black holes, known as the gravitational vacuum condensate star, or gravastar [36], which was further elaborated in 2004 [37]. Related ideas were independently explored by Chapline and collaborators [38, 39, 40], who argued that quantum effects could render the black hole horizon a critical surface associated with a gravitational phase transition. In this picture, the interior of the object is described by an equation of state (EOS) of the form $p=-\rho$ [41], corresponding to a vacuum energy that exerts a repulsive gravitational effect capable of counteracting collapse.
@@
Building on these ideas, MM suggested that quantum fluctuations become dominant near the would-be horizon, particularly in the temporal and radial components of the energy–momentum tensor. These fluctuations can grow sufficiently large to effectively produce an EOS $p=\rho$, which represents the stiffest causal equation of state allowed by relativity. Such a state lies at the threshold of causality violation and naturally leads to the formation of a gravitational Bose–Einstein condensate (BEC) in the interior. The sharp horizon is then replaced by a thin shell composed of stiff matter, a concept first introduced by Zeldovich [42, 43]. Outside this shell, the spacetime is assumed to be vacuum, with vanishing pressure and energy density.
The gravastar model thus emerges as a non-singular endpoint of gravitational collapse, designed specifically to eliminate the central curvature singularity that plagues classical black hole solutions [36]. In the simplest Schwarzschild black hole scenario, the event horizon is a null hypersurface across which physical quantities remain finite, allowing infalling observers to cross it smoothly. However, whether such analytic continuation remains valid once quantum effects are properly taken into account remains an open question. Motivated by this uncertainty, MM replaced the classical event horizon with a thin shell separating distinct interior and exterior phases [37].
@@
In the standard MM construction, the interior region is modeled as a vacuum condensate characterized by the EOS $p=-\rho$, which is mathematically equivalent to a cosmological constant and generates a repulsive gravitational effect. Although this EOS violates the strong energy condition (SEC), $\rho+3p\geq 0$, it still satisfies the null energy condition (NEC), $p+\rho\geq 0$. Since violation of the NEC is a key ingredient in the Hawking–Penrose singularity theorems [6], it is not immediately obvious that this alone is sufficient to prevent singularity formation. Nevertheless, the violation of the SEC introduces repulsive effects that can significantly alter the collapse dynamics. To incorporate quantum considerations more fully, following Chapline’s arguments [38, 40], the horizon is treated as a locus of quantum phase transition, supported by an interior that violates the SEC and allows for a rearrangement of the vacuum state [41]. The resulting configuration consists of a gravitational BEC core stabilized by an ultra-dense thin shell of stiff matter obeying $p=\rho$ [42, 43]. Comprehensive reviews of the gravastar paradigm can be found in [44].
@@
In recent years, gravastars have attracted considerable interest, and a variety of stable configurations have been constructed within both GR and modified gravity theories [45, 46, 47, 48, 49, 50, 51]. Since the avoidance of singularities is fundamentally a UV effect, it is particularly compelling to investigate gravastar solutions in frameworks that modify GR at high energies, where quantum gravitational corrections are expected to play a dominant role. The motivation for considering gravastars in a braneworld framework arises from the ultraviolet sensitivity of gravitational collapse, where classical four-dimensional general relativity is expected to receive significant corrections. While gravastar solutions can be constructed in GR, their stability typically requires imposed anisotropies, exotic equations of state, or infinitesimally thin shells. In braneworld gravity, the effective four-dimensional field equations acquire high-energy quadratic stress–energy corrections and non-local Weyl terms sourced by the bulk geometry, which become dominant at the energy scales relevant for compact objects. These geometrically induced contributions naturally generate effective repulsive gravity, intrinsic pressure anisotropy, and a reduced (or negative) active gravitational mass on the brane, thereby providing a theoretically grounded mechanism for stabilizing non-singular gravastars without introducing ad hoc matter sectors, while simultaneously offering potentially observable deviations from classical black hole phenomenology.
@@
This work is devoted to the construction and analysis of a gravastar configuration within the framework of the SS braneworld scenario. The paper is structured as follows. In Sec. 2, we present the effective field equations on a SS brane for a static, spherically symmetric spacetime and derive their explicit solutions by separately imposing appropriate equations of state for the interior region, the thin shell, and the exterior spacetime of the gravastar. A detailed examination of the resulting physical characteristics of the gravastar model is carried out in Sec. 3. The matching conditions at the interfaces between different regions are established in Sec. 4. Subsequently, the dynamical stability of the proposed gravastar configuration is analyzed in Sec. 5. Finally, a summary of the main findings along with concluding remarks is provided in Sec. 6.
2 Mathematical formalism and solutions

### v2

# Gravastars on the brane with a timelike extra dimension

@@ Abstract
We construct a gravastar configuration within the Shtanov-Sahni (SS) braneworld scenario with a timelike extra dimension and negative brane tension. The modified gravitational dynamics on the brane naturally regularize the interior geometry and prevent central curvature singularities. Solving the induced field equations, we obtain analytic interior, shell, and exterior solutions without invoking the thin-shell approximation. The gravastar core is modeled as a Bose–Einstein condensate, while the intermediate shell consists of ultra-dense stiff matter. Bulk Weyl corrections generate anisotropic effective pressures that arise intrinsically and support stability. We analyze the active gravitational mass, energy, entropy, proper shell thickness, and junction conditions at the interfaces. The configuration can exhibit suppressed or negative effective mass due to the repulsive interior condensate while remaining consistent with energy conditions, demonstrating that SS braneworld gravastars provide a viable compact-object alternative to classical black holes with singularity avoidance arising from higher-dimensional geometric effects.
I Introduction
@@
The question of how gravitational collapse ultimately ends in astrophysics mirrors, in many respects, the problem of the initial state of the universe in cosmology. While Einstein’s General Relativity (GR) has enjoyed remarkable success in explaining a wide range of observational phenomena at intermediate energy scales in both astrophysical and cosmological contexts, its predictive power is severely challenged in regimes of extreme curvature and energy density. Such regimes naturally arise at the putative beginning of the universe and at the final stages of stellar gravitational collapse. In these limits, enormous amounts of energy are compressed into extremely small volumes, causing the energy density and curvature invariants to diverge. As a consequence, GR generically predicts spacetime singularities where the classical field equations lose their validity Hawking1 . This pathological behavior strongly suggests that new physical effects, most plausibly of quantum origin, must become relevant in order to regularize the theory and remove these unphysical infinities.
@@
In the context of stellar collapse, incorporating quantum field effects on a fixed classical spacetime background has already revealed several profound phenomena. The most celebrated example is Hawking radiation, whereby black holes emit thermal radiation from their event horizons Hawking2 , often interpreted as the boundary separating the black hole interior from the external universe. Despite this remarkable insight, semiclassical gravity is insufficient to cure the singular behavior of classical black hole solutions, as the curvature singularity at the center persists unchanged Birrell . In the Schwarzschild solution, the apparent singularity at the Schwarzschild radius R=2GM(c=1)R=2GM~(c=1) is merely a coordinate artifact, since all curvature invariants remain finite there and the singularity can be eliminated by an appropriate change of coordinates. In contrast, the divergence at $r=0$ corresponds to a genuine physical singularity, characterized by divergent curvature scalars, and cannot be removed by any coordinate transformation.
@@
From a fundamental perspective, the emergence of physical singularities in the description of any physical system is widely regarded as signaling the breakdown or incompleteness of the underlying theoretical framework. Relativistic gravity presents two prominent and closely related examples of such singularities: those associated with cosmological models of the universe Hawking and those arising in black hole spacetimes Penrose . Both are curvature singularities, meaning that they are invariantly defined through divergences of curvature tensors and cannot be eliminated by coordinate redefinitions HP . In standard relativistic cosmology, a Ricci-type curvature singularity occurs at $t=0$, where the Riemann tensor and the energy density diverge, marking what is conventionally interpreted as the beginning of spacetime evolution Hawking . Analogously, classical black hole solutions contain a curvature singularity at $r=0$, where physical quantities blow up in an identical manner Penrose . This central singularity stands in sharp contrast to the horizon at $r=2M$, which represents only a removable coordinate singularity in Schwarzschild coordinates and can be regularized through alternative coordinate systems  Finkelstein ; Kruskal ; Szekeres .
@@
In modern theoretical physics, it is increasingly accepted that the presence of such singularities reflects the limited domain of applicability of classical GR rather than a true physical feature of nature. This viewpoint is motivated by two key considerations. First, GR has not been experimentally tested in regimes of ultra-high curvature or energy density. Second, it is natural to expect that quantum effects should dominate the dynamics at such extreme scales, whereas GR is intrinsically a classical theory. Consequently, substantial effort has been devoted to extending or modifying GR in a manner that remains consistent with known observations while improving its high-energy behavior. One particularly influential strategy is to modify the geometric sector of the theory by generalizing the Einstein–Hilbert (EH) action. Originally introduced several decades ago as a means of incorporating ultraviolet (UV) corrections, such modifications have since been widely explored at both UV and infrared (IR) scales. This approach is often favored because it avoids many of the conceptual and phenomenological difficulties that arise when the matter sector alone is altered. Moreover, modifying the gravitational action may provide valuable clues toward the eventual construction of a self-consistent theory of quantum gravity (QG), even though such a complete theory remains elusive at present. In this regard, the early universe and compact objects like black holes constitute natural laboratories for probing the viability of QG-inspired models, as both are characterized by extreme curvature and energy density. Among the various approaches to quantum gravity, two broad and conceptually distinct frameworks have gained widespread attention. One of these is M-theory, which emerged as a unifying framework linking the five previously known superstring theories through duality relations and which reduces to supergravity in the low-energy limit S2 . Quantum consistency of M-theory requires an eleven-dimensional spacetime, and its fundamental degrees of freedom include extended objects such as two- and five-dimensional membranes, commonly referred to as branes. Although M-theory has not yet been fully formulated, its structural features have inspired numerous developments in high-energy physics and cosmology.
@@
Braneworld models have emerged as an important extension of higher-dimensional gravity, with the Randall–Sundrum (RS) Randall1 ; Randall2 playing a central role. This scenario provides an elegant explanation of the hierarchy between the gravitational and electroweak scales by postulating that our observable universe is a $(3+1)$-dimensional hypersurface embedded in a higher-dimensional spacetime. In this setup, Standard Model fields are confined to the brane, whereas gravity propagates in the surrounding five-dimensional bulk endowed with an Anti–de Sitter geometry ($AdS_{5}$). The presence of a negative bulk cosmological constant leads to the localization of the graviton near the brane. While the original RS proposal involved two branes, the later RS-II formulation Randall2 effectively removes the negative-tension brane by sending it to infinity, yielding a single-brane configuration. The Randall–Sundrum single–brane framework with a spacelike extra dimension has been widely investigated and applied to a variety of problems in cosmology Binetruy ; Maeda ; Langlois ; Chen ; Kiritsis ; Campos ; Maartens as well as in astrophysical contexts Wiseman2 ; Germani ; Wiseman ; Visser ; Creek ; Pal ; Bruni ; Govender ; Sengupta5 . Building on this line of research, Shtanov and Sahni proposed an alternative construction, commonly referred to as the SS model, which may be regarded as a dual of the RS–II scenario, where the spacelike extra dimension is replaced by a timelike one Sahni4 . In the standard RS case the five–dimensional bulk possesses a Lorentzian signature, whereas introducing a timelike extra dimension leads to a bulk geometry with signature (,,+,+,+)(-,-,+,+,+). This change in the nature of the extra dimension significantly modifies the properties of the braneworld. Unlike the RS–II model, the SS setup is characterized by a negative brane tension and a positive bulk cosmological constant. One of its most remarkable features is that the cosmological evolution admits a smooth bounce during the contracting phase, thereby avoiding the formation of a singularity, even when the matter content on the brane satisfies the usual energy conditions. A potential concern in models with a timelike extra dimension is the emergence of tachyonic Kaluza–Klein gravitational modes, as originally discussed in Ref. Sahni4 . This issue has, however, been subsequently examined and addressed in Ref. S10 .
@@
Against this broader backdrop, a novel alternative to classical black holes was proposed in the form of the gravitational vacuum condensate star, or gravastar Mazur1 , which was further developed in subsequent work Mazur2 . In parallel, related ideas suggested that quantum gravitational effects could transform the classical black hole horizon into a critical surface associated with a gravitational phase transition. In this picture, the interior of the object is described by an equation of state (EoS) of the form $p=-\rho$ Gliner , corresponding to a vacuum energy that exerts a repulsive gravitational effect capable of counteracting collapse. Building on these ideas, it was suggested that quantum fluctuations become dominant near the would-be horizon, particularly in the temporal and radial components of the energy–momentum tensor. These fluctuations can grow sufficiently large to effectively produce an EoS $p=\rho$, which represents the stiffest causal equation of state allowed by relativity. Such a state lies at the threshold of causality violation and naturally leads to the formation of a gravitational Bose–Einstein condensate (BEC) in the interior. The sharp horizon is then replaced by a thin shell composed of stiff matter, a concept first introduced by Zeldovich Zeldo1 ; Zeldo2 . Outside this shell, the spacetime is assumed to be vacuum, with vanishing pressure and energy density. Early seminal works established primordial black holes (PBHs) as natural outcomes of scalar-field instabilities and first-order cosmological phase transitions, linking PBH formation to symmetry-breaking dynamics in the early universe Khlopov1985 ; Konoplich1999 ; Khlopov2000 . These studies demonstrated how bubble collisions, false-vacuum collapse, and density inhomogeneities can efficiently seed black holes across a wide mass spectrum. Subsequent reviews consolidated PBHs as probes of high-energy physics and cosmology beyond the Standard Model Khlopov2010 . More recently, regular black hole remnants with de Sitter cores and associated graviatoms were proposed as viable heavy dark matter candidates, sensitive to primordial inhomogeneities Dymnikova2015 .
@@
The gravastar model thus emerges as a non-singular endpoint of gravitational collapse, designed specifically to eliminate the central curvature singularity that plagues classical black hole solutions Mazur1 . In the simplest Schwarzschild black hole scenario, the event horizon is a null hypersurface across which physical quantities remain finite, allowing infalling observers to cross it smoothly. However, whether such analytic continuation remains valid once quantum effects are properly taken into account remains an open question. Motivated by this uncertainty, the classical event horizon was replaced by a thin shell separating distinct interior and exterior phases Mazur2 . In the standard gravastar construction, the interior region is modeled as a vacuum condensate characterized by the equation of state ($p=-\rho$), which is mathematically equivalent to a cosmological constant and generates a repulsive gravitational effect. Although this EOS violates the strong energy condition (SEC), $\rho+3p\geq 0$, it still satisfies the null energy condition (NEC), $p+\rho\geq 0$. Since violation of the NEC is a key ingredient in the Hawking–Penrose singularity theorems HP , it is not immediately obvious that this alone is sufficient to prevent singularity formation. Nevertheless, the violation of the SEC introduces repulsive effects that can significantly alter the collapse dynamics. To incorporate quantum considerations more fully, the horizon is treated as a locus of quantum phase transition, supported by an interior that violates the SEC and allows for a rearrangement of the vacuum state Gliner . The resulting configuration consists of a gravitational BEC core stabilized by an ultra-dense thin shell of stiff matter obeying $p=\rho$ Zeldo1 ; Zeldo2 . Comprehensive reviews of the gravastar paradigm can be found in Ref. RS1 .
In recent years, gravastars have attracted considerable interest, and a variety of stable configurations have been constructed within both GR and modified gravity theories G1 ; G2 ; G4 ; G5 ; G6 ; Sengupta5 . Such investigations are particularly relevant in the strong-field regime, where deviations from general relativity may become significant and potentially observable. The three-dimensional gravastar structures in massive gravity Barzegar:2023plb and in gravity’s rainbow Barzegar:2023epjc provide novel insights into lower-dimensional models, while extensions in $f(Q)$ gravity have been systematically explored through Chaplygin gas models Mohanty:2025ijmpd , Krori–Barua metrics Mohanty:2024fdp , and charged configurations Mohanty:2024aop . Charged gravastars in $f(R,T)$ gravity have also been studied in detail Yousaf:2019prd . A recent work explored gravastar configurations in $f(\mathcal{G},T)$ gravity, showing that Gauss–Bonnet curvature and matter–geometry couplings can support regular, stable horizonless compact objects with properties distinct from general relativity Shamir2018 . More recently, another work studied thin-shell gravastars in massive gravity using the Darmois–Israel formalism, demonstrating that a non-zero graviton mass significantly modifies shell stresses and stability conditions. Their results indicate that both higher-curvature corrections and massive gravity provide viable mechanisms for realizing stable, non-singular gravastar configurations without event horizons Das2024 .
Recent studies have significantly advanced the understanding of gravastars as viable horizonless compact objects across dynamical, perturbative, and observational contexts. It has been showed that gravastar formation via sudden vacuum condensation can be accompanied by thermal radiation with a characteristic de Sitter temperature, providing a possible radiative signature of gravastar birth Nakao2023 . Adler demonstrated that time-dependent gravastar configurations can evade several no-go results applicable to static exotic compact objects, underscoring the importance of dynamics for stability and light propagation Adler2024 . Some works have investigated observational imprints of thin-shell gravastars using ray-tracing simulations of accretion disks and hot spots, finding that gravastars can closely mimic black holes while still exhibiting potentially detectable deviations Rosa2024 . From a perturbative perspective, analyzed scalar quasinormal modes of gravastars have been analyzed in scalar–tensor theories, showing that additional scalar degrees of freedom modify the mode spectrum relative to general relativity Boumaza2025 . Earlier,it was established that gravastar oscillation spectra and gravitational-wave signatures differ qualitatively from those of black holes Pani2009 , while others have demonstrated that gravastars can approach the black-hole limit in their tidal deformability, clarifying how I–Love–Q relations for gravastars connect smoothly to those of black holes Uchikata2016 .
@@
Since the avoidance of singularities is fundamentally a UV effect, it is particularly compelling to investigate gravastar solutions in frameworks that modify GR at high energies, where quantum gravitational corrections are expected to play a dominant role. The motivation for considering gravastars in a braneworld framework arises from the ultraviolet sensitivity of gravitational collapse, where classical four-dimensional general relativity is expected to receive significant corrections. While gravastar solutions can be constructed in GR, their stability typically requires imposed anisotropies, exotic equations of state, or infinitesimally thin shells. In braneworld gravity, the effective four-dimensional field equations acquire high-energy quadratic stress–energy corrections and non-local Weyl terms sourced by the bulk geometry, which become dominant at the energy scales relevant for compact objects. These geometrically induced contributions naturally generate effective repulsive gravity, intrinsic pressure anisotropy, and a reduced (or negative) active gravitational mass on the brane, thereby providing a theoretically grounded mechanism for stabilizing non-singular gravastars without introducing ad hoc matter sectors, while simultaneously offering potentially observable deviations from classical black hole phenomenology.
@@
This work is devoted to the construction and analysis of a gravastar configuration within the framework of the SS braneworld scenario. The paper is structured as follows. In Sec. II, we present the effective field equations on a SS brane for a static, spherically symmetric spacetime and derive their explicit solutions by separately imposing appropriate equations of state for the interior region, the thin shell, and the exterior spacetime of the gravastar. A detailed examination of the resulting physical characteristics of the gravastar model is carried out in Sec. III. The matching conditions at the interfaces between different regions are established in Sec. IV. Subsequently, the dynamical stability of the proposed gravastar configuration is analyzed in Sec. V. Finally, a summary of the main findings along with concluding remarks is provided in Sec. VI.
II Mathematical formalism and solutions